OpenStudy (anonymous):

Write an equation of the cosine function with amplitude 2 and period 8π.

OpenStudy (anonymous):

f(x)=2cos(x/4) Let's verify if written function meets the requirements: |2cos(x/4)|<=2 -amplitude f(x+8pi)=2cos((x+8pi)/4)=2cos(x/4+2pi)=2cos(x/4)=f(x) - so 8pi is the period